OpenStudy (anonymous):

*physics question(constant acceleration)*At a stoplight, a truck traveling at 15 m/s passes a car as it starts from rest.?

OpenStudy (anonymous):

wrong board and where is the rest of the question?

OpenStudy (anonymous):

At a stoplight, a truck traveling at 15 m/s passes a car as it starts from rest. The truck travels at constant velocity and the car accelerates at 3 m/s^2. How much time does the car take to catch up to the truck?

OpenStudy (anonymous):

well, you just have two equations to set equal to one another

OpenStudy (anonymous):

d=v*t+1/2*a*t^2 is the general form of these equations i believe

OpenStudy (anonymous):

so 15*t=1/2*3*t^2

OpenStudy (anonymous):

so 10*t=t^2

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

therefore, t=10 or 0 and since 0 is when we start, t=10

OpenStudy (anonymous):

how did you get 15? i know its right but i dont know how to get there

OpenStudy (anonymous):

the truck is traveling at 15 m/s, so its distance equation is d=15*t

OpenStudy (anonymous):

the car is starting from rest (v=0) and accelerating at 3m/s. so its distance equation is d=1/2*3*t^2

OpenStudy (anonymous):

i just set those two equal to one another and solved

OpenStudy (anonymous):

oh i see thanks, a lot i know this isnt the right place to ask questions but more people are here

OpenStudy (anonymous):

lol, kk